OpenStudy (anonymous):

IT'S ALMOST STRUCK MIDNIGHT !!! HELP !!!! Solve the inequality. 5x - 6 (>=) 9 A) x (>=) 5 B) x > -3 C) x (>=) 3 D) x (<=) 5

OpenStudy (shinalcantara):

for you to find the value of 'x' you'll have it as: \[5x - 6 \ge 9\] add 6 to both sides of the inequality \[5x - 6 + 6 \ge 9 + 6\] \[5x \ge 15\] divide the whole inequality with 5 \[x \ge 3\]

OpenStudy (anonymous):

THANK YOU !!!